Operations Report – Feb 27th

Crew 222 Operations Report 27FEB2020

SOL: 11

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Wednesday generator:

Off at 6:00am

On at 9:00pm SOC was 85%

Thursday generator:

Off at 5:30am

On at 9:00pm SOC was 83%

Diesel Reading – 72%

Station Propane Reading – 84%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 200 gallons

Water Meter: 149548.32 units

Water (static tank) – 218 gallons

Static to Loft Pump used – yes

Water in GreenHab – 35.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes

Hours: 146.7

Beginning charge: 100%

Ending charge: 89%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Spirit rover: no

Hours: 132.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Curiosity rover used: no

Hours: 138.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1 and 2)

Reason for use: Run during the engineer EVA for the battery and for the

EVA this morning.

Oil Added? no

ATV Fuel Used: 20% of the tank.

Hours the ATVs were Used today: 30 minutes

Notes on ATVs: Oil checked. The Honda needs an oil change. The four ATVs

were operational this afternon.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.

The suit number 7 was working well at the beginning but two hours after the beginning of the EVA the fan was stopping at random moments but was restarting after. The fan never stopped totally but wasn’t working optimally.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report